3x to the 2/5 - 5 = 22
Answer by stanbon(57347)

3x to the 2/5 - 5 = 22
--------------
(3x)^(2/5) - 5 = 22
(3x)^(2/5) = 27
(3x)^(2/5) = 3^3
Raise both sides to the (5/2) power
3x = 3^(3*(5/2))
3x = 3^(15/2)
x = 3^[(15/2)-1]
x = 3^[13/2]
x = 1262.67
===================================
Cheers,
Stan H.